OpenStudy (anonymous):

if angle ABC =5 and angle 2 = 40 then what is angle 1?

OpenStudy (anonymous):

135 degrees

OpenStudy (anonymous):

ABC = 125 and angle 2 = 40 what is angle 1

OpenStudy (anonymous):

180-(125+40) = 180- 165 = 15 degrees